OpenStudy (yttrium):

A ball is dropped from a boat so that it strikes the surface of a lake with a speed of 16.5 fps. While in the water the ball experiences an acceleration of a=10−0.8v, where a and v are expressed in fps^2 and fps respectively. Knowing the ball takes 3 sec to reach the bottom of the lake, determine the following: a. the depth of the lake b. the speed of the ball when it hits the bottom of the lake

OpenStudy (yttrium):

@ganeshie8

OpenStudy (yttrium):

@satellite73

OpenStudy (anonymous):

fps = feet per sec?

OpenStudy (yttrium):

yes

OpenStudy (anonymous):

Is \(a=10-0.8v\) a function of a variable velocity \(v\)? If so, you can try solving for \(v\) since this is a differential equation.

OpenStudy (anonymous):

can one write ball will float and never go down to lake's surface?

OpenStudy (yttrium):

yes it is.

OpenStudy (yttrium):

can you help me @SithsAndGiggles

OpenStudy (anonymous):

Do you know how to solve linear DEs? Find the integrating factor, etc? I'm not sure what to do beyond that, but you'll have a function of the velocity with respect to time. That along with whatever formulas you've learned should be enough to figure out what you need. I don't know if you treat it like a free fall problem or not, though...

OpenStudy (yttrium):

i know how to solve linear DEs. But I don't know how to deal with it.

OpenStudy (yttrium):

I mean with this problem

OpenStudy (anonymous):

Use SUVAT equations :)

OpenStudy (anonymous):

You're given an initial velocity of \(v(0)=16.5\), so the initial acceleration is \(a(16.5)=10-0.8(16.5)=-3.2\). Let's see what we get for the velocity function: \[\large\begin{align*}\frac{dv}{dt}+0.8v&=10\\ e^{0.8t}\frac{dv}{dt}+0.8e^{0.8t}v&=10e^{0.8t}\\ \frac{d}{dt}\bigg[e^{0.8t}v\bigg]&=10e^{0.8t}\\ v&=10e^{-0.8t}\left(\int e^{0.8t}~dt\right)\\ v&=10e^{-0.8t}\left(\frac{1}{0.8}e^{0.8t}+C\right)\\ v&=12.5+Ce^{-0.8t}\\ 16.5&=12.5+C\\ C&=4 \end{align*}\] So \(v(t)=12.5+4e^{-0.8t}\)... The ball takes three second to reach the bottom. Presumably, the ball stops moving at that time (doesn't bounce back up), so (I would think) you have velocity defined by \[v(t)=\begin{cases}12.5+4e^{-0.8t}&\text{for }0\le t<3\\0&\text{for }t\ge3\end{cases}\] This makes me think that plugging in \(t=3\) should give the velocity of the ball as it hits the bottom, which would be about 12.86 fps. Not sure how to approach part (a) just yet...

OpenStudy (anonymous):

Actually, I just might... I think it's just a matter of integrating the velocity function over the interval [0,3], which would give the total displacement. Since the ball is traveling straight down (with no obstructions), the depth of the water would be the displacement. \[d=\int_0^3 v(t)~dt\approx42.05\]

OpenStudy (phi):

sith's solution looks good. you could also solve using a= x'' , v= x' i.e. distance with respect to time you get the 2nd order diff eq x'' +0.8 x' = 10 solve the homogeneous equation x'' + 0.8 x' = 0 which has the characteristic equation m^2 + 0.8 m =0 m=0 and m= -0.8 thus \[ x_h= c_1 e^{0t} + c_2 e^{-0.8t} \\ x_h= c_1 + c_2 e^{-0.8t} \] for the particular solution, we would choose a constant x=C except that the homogeneous sol'n already has a constant. so instead use xp= Ct putting that into the equation we get 0.8C = 10, C= 10/.8 = 12.5 the solution for distance traveled is \[ x= x_h+x_p = c_1 + c_2 e^{-0.8t} + 12.5t \] to find the constants, we use x'= 16.5 at t=0 differentiate x to find \[ x'= -0.8 c_2 e^{-0.8\cdot 0} +12.5 = 16.5 \] we find \( c_2= - 5 \) we set x=0 at t=0 and solve \[ c_1 - 5 e^{-0.8\cdot 0} + 12.5 \cdot 0 =0 \\ c_1= 5\] and finally \[x= 5 - 5 e^{-0.8t} + 12.5t \] and velocity \[ x'= v = 4 e^{-0.8t} + 12.5 \] evaluate at t=3

OpenStudy (yttrium):

what is \[x_{h}\] by the way?

OpenStudy (phi):

the solution of a differential equation has two parts: 1) the particular solution. in this case, the solution that gives 10 2) the homogeneous solution, where we equate the equation to 0 rather than to 10. See http://tutorial.math.lamar.edu/Classes/DE/HOHomogeneousDE.aspx for a discussion.

OpenStudy (yttrium):

Ok, you mean Xh to be the characteristic solution?

OpenStudy (yttrium):

can you just explain briefly where 0.8C = 10 came from?

OpenStudy (phi):

To find the particular solution, we "guess" it. if we have equation = 10 we would guess the particular solution is 10 and all possible derivatives. But if the homogeneous solution already has a constant as part of its answer (and in your problem it does) we try x= C t (we multiply by the independent variable) as our guess. see http://tutorial.math.lamar.edu/Classes/DE/UndeterminedCoefficients.aspx we put our guess into the diff equation x'' + 0.8 x' = 10 with x= C t, we find x' = C (i.e. d/dt C t = C) and x'' = d/dt C = 0 we get 0 + 0.8 C = 10 or 0.8 C =10 solve for C to get 12.5 and the particular solution xp= 12.5 t

OpenStudy (yttrium):

where does x = Ct equation came? Is that just a guess?

OpenStudy (phi):

yes, that is the guess. Paul's discussion might help (though I have not read through it, he's generally helpful)

OpenStudy (yttrium):

but where does that guess came from? how do you arrive at that guess?

OpenStudy (phi):

This is the "method of undetermined coefficients" and it only works for when the right hand side is a sum of terms, each term having a finite number of linearly independent derivatives. This restricts the technique to terms such as \[ c, t^c, e^{ct}, \sin c\ t, \cos c\ t \] luckily we can use it for constant right-hand side (i.e. the 10)